OpenStudy (vane11):

Conversion Question? I have .00199888g/L and I'm supposed to convert to L/kg The TA gave us this to follow: g/kg x L/g = L/kg Can someone do it for me to see if it checks out with mine? I suck at simple conversions

OpenStudy (nincompoop):

that just means grams per liter let me give you the first step: 0.00199888 g/L x (1kg/1000g)

OpenStudy (nincompoop):

your g cancels out and you are left with Kg, now just do the arithmetic

OpenStudy (nincompoop):

you have to do the math... I am merely guiding you

mathslover (mathslover):

\(0.00199888 \cancel{g} / L \times \cfrac{1 \space kg}{1000 \cancel{g}}\) ^ Just texed.. @nincompoop 's work.

OpenStudy (nincompoop):

yo how do you do that cancel thing? it looks awesome LOL

OpenStudy (vane11):

it does.... lol and so I end up with kg/L right? Well with the numbers in front of it

OpenStudy (nincompoop):

so what do you have?

mathslover (mathslover):

\cancel

OpenStudy (nincompoop):

metric system was invented for a reason... that is to make conversions easier

OpenStudy (vane11):

1.999x10^-6kg/L

mathslover (mathslover):

\cancel{_} `\[\cancel{g} \times \cfrac{kg}{\cancel{g}}\] `

OpenStudy (nincompoop):

you're dividing by 1000 that means your decimal place moves 3 steps to the left and don't leave out the rest of the remaining digits

OpenStudy (nincompoop):

so what do you think is next based on what your TA gave you?

OpenStudy (vane11):

yeah that's why I multiplied it by 10 to the negative 6 power and I think he may have meant kg/L because next I have to multiply it by my weight in kilograms to find out how many Liters of this substance it would take to kill me... haha I know it sounds wierd

OpenStudy (vane11):

does this look right? 1.999x10^-6kg/L x 106lbs/48.08kg x 1lb

OpenStudy (vane11):

Ok thanks didn't know that...

OpenStudy (nincompoop):

using the formula: Density = mass/volume your format is going to be Kg/L 1.999888 x 10^-6 Kg/L

OpenStudy (nincompoop):

to pounds? I didn't see that in the original problem ...

OpenStudy (vane11):

Well I know my weight in pounds so I converted it to kg

OpenStudy (vane11):

106 x .453592 = 48.080752

OpenStudy (nincompoop):

lmao! you are right, but you can leave out x lb.

OpenStudy (vane11):

yeah since it's 1 haha

OpenStudy (nincompoop):

no, because it's like this \[Rabbit \times \frac{ Frog }{ Rabbit }=Frog\]

OpenStudy (vane11):

well I knew what I meant... haha

OpenStudy (vane11):

thanks for helping me out it's for my lab report :)

OpenStudy (chmvijay):

LOL:)

OpenStudy (vane11):

my baaad haha and sure